Evaluate the moment of inertia of a thin rod with axis of rotation at the centre , if its length is 100cm and its mass is 2.3kg.


1
Expert's answer
2022-01-23T16:13:53-0500

"I=\frac{mL^2}{12}=\frac{2.3\cdot 1^2}{12}=0.19~kg\cdot m^2."